OpenStudy (anonymous):

can someone teach me how to determine if a line is para or perp or neither l1 y=x-3 l2 y=1-x

jimthompson5910 (jim_thompson5910):

What is the slope of each line

OpenStudy (anonymous):

for slope is it y=mx+b

jimthompson5910 (jim_thompson5910):

that's the slope intercept form m is your slope

OpenStudy (anonymous):

i dont get it, so was looking for examples on how to set the problem up

jimthompson5910 (jim_thompson5910):

y=x-3 is the same as y=1x-3

jimthompson5910 (jim_thompson5910):

y = 1x - 3 is in the form y = mx+b where m = 1, b = -3

jimthompson5910 (jim_thompson5910):

so the slope of y = x-3 is 1

jimthompson5910 (jim_thompson5910):

what's the slope of y = 1 - x

OpenStudy (anonymous):

what happens when x is not the like for line 2 will it still be 1?

jimthompson5910 (jim_thompson5910):

well y = 1 - x is the same as y = -x + 1 which is the same as y = -1x + 1

OpenStudy (anonymous):

is that a rule?

jimthompson5910 (jim_thompson5910):

the slope of this is what?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

once you set it up like this do we do m=y2-x2/y1-x2

jimthompson5910 (jim_thompson5910):

no need for that since we can easily pick out the slope

OpenStudy (anonymous):

the m is the slope so its 1 and -1?

jimthompson5910 (jim_thompson5910):

are those slopes the same

OpenStudy (anonymous):

no

OpenStudy (anonymous):

para would be same and perp will be different?

jimthompson5910 (jim_thompson5910):

parallel lines would have equal slopes

jimthompson5910 (jim_thompson5910):

perpendicular lines would have 2 slopes that multiply to -1

jimthompson5910 (jim_thompson5910):

so 1 times -1 = -1 which means that these two lines are perpendicular

OpenStudy (anonymous):

what would make it neither?

OpenStudy (anonymous):

thanks for explaining it by the way

jimthompson5910 (jim_thompson5910):

if the slopes were different AND they didn't multiply out to -1